OpenStudy (clamin):

Find the value of c that completes the square. p^2-20p+c

TheSmartOne (thesmartone):

(-20/2)^2 -10^2 100=c

OpenStudy (dumbcow):

in general, you have; \[p^2 +bp + c = (p+ \frac{b}{2})^2\] so \[c = (\frac{b}{2})^2\] here b = -20